OpenStudy (anonymous):

Find the LA of a regular hexagonal pyramid w/ e=20m and the radius base is 9sqrt of 3 cm.

OpenStudy (michele_laino):

here the side of the hexagon is: \[\frac{{\sqrt 3 }}{2}l = 9\sqrt 3 \] so l= 18 meters

OpenStudy (anonymous):

@CGGURUMANJUNATH